# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v1

@@ 5.1 Linear and nonlinear univariate time series
Figure 3: Scatter plots of $(x_{t-h},x_{t})$ of the original data and $(\hat{w}_{t-h},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 5 (CYC) under missing pattern I. $h\in\{1,2,6,10\}$.
@@
Next, we examine the sample autocovariance functions (ACFs) estimated from the imputed data. Table 2 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$. For Model 1 (AR), most methods perform reasonably well. Linear interpolation produces accurate ACF estimates since the positive AR coefficient intrinsically favors linearity, while the scalar filter exhibits largest estimation errors. For Model 2 (ARMA), TWI improves from linear and Kalman initializations and yields more accurate ACF estimates: TWI reduces the biases in Kalman smoothing when $h=0$ and in linear interpolation when $h=1$, and $h=2$. For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series exhibit the lowest RMSEs, with substantial improvements over the initializations. Notably, the scalar filter also yields accurate ACFs for $h=1,2$. Figure 4 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
@@
Figure 4: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR). The blue line is the smoothed density estimated from the histogram of the ACFs estimated from the ground truth.
Table 2: $\mathrm{RMSE}_{\gamma}(h)$, defined in (19), for $h=0,1,2$, with $h$ denoting lag. The results are averaged over 1000 simulations, and the smallest values in each row are marked in boldface. See Table 1 for notations.
@@
Finally, we turn to the model parameters estimated from the imputed series. Table 3 presents the RMSEs for each model parameter. For Model 1 (AR) and Model 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements. For the nonlinear Model 3 (TAR), the results are drastically different. TWI produces the most accurate parameter estimates and can achieve approximately 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing. In some cases, such as with linear initialization under missing pattern II, the improvement far exceeds 50%. These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics. For Model 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization. Figure 5 shows the histogram of the estimated parameter $\phi_{1}$ in Model 4. It is visible from the figure that TWI helps neutralize the biases from the initialization and produces reliable estimates.
Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
Missing pattern I
AR $\phi$ 0.06 0.06 0.05 0.03 0.03 0.04 0.03 0.04
ARMA $\phi_{1}$ 0.19 0.47 0.12 0.08 0.17 0.16 0.11 0.11
$\phi_{2}$ 0.46 0.95 0.18 0.10 0.33 0.26 0.16 0.15
TAR $\phi_{1}$ 1.02 1.67 0.82 0.63 0.71 0.58 0.53 0.44
$\phi_{2}$ 0.08 0.07 0.20 0.04 0.05 0.03 0.03 0.02
$\tau$ 0.01 0.75 0.35 0.20 0.05 0.05 0.03 0.03
I(1) $\phi$ 0.96 0.47 0.45 0.68 0.48 0.32 0.30 0.24
NLVAR $\phi_{11}$ 0.19 0.27 0.15 0.03 0.06 0.03 0.05 0.03
$\phi_{12}$ 3.10 3.16 2.87 1.67 1.34 1.02 1.22 0.97
$\phi_{21}$ 0.03 0.07 0.03 0.03 0.03 0.03 0.03 0.03
$\phi_{21}$ 0.18 0.23 0.10 0.04 0.05 0.04 0.04 0.04
Missing pattern II
AR $\phi$ 0.05 0.07 0.04 0.04 0.03 0.02 0.03 0.02
ARMA $\phi_{1}$ 0.04 0.08 0.09 0.08 0.05 0.06 0.09 0.09
$\phi_{2}$ 0.16 0.61 0.11 0.10 0.12 0.11 0.11 0.12
TAR $\phi_{1}$ 0.72 2.44 0.38 0.25 0.25 0.20 0.29 0.21
$\phi_{2}$ 0.09 0.12 0.05 0.03 0.06 0.05 0.02 0.01
$\tau$ 0.01 1.72 0.01 0.01 0.01 0.01 0.01 0.01
I(1) $\phi$ 0.08 1.08 0.11 0.70 0.09 0.10 0.10 0.10
NLVAR $\phi_{11}$ 0.19 0.44 0.08 0.02 0.02 0.01 0.02 0.01
$\phi_{12}$ 2.67 3.51 1.68 1.26 0.61 0.46 0.61 0.43
$\phi_{21}$ 0.02 0.09 0.03 0.03 0.06 0.05 0.04 0.03
$\phi_{21}$ 0.15 0.34 0.05 0.04 0.04 0.04 0.04 0.04
(a) missing pattern I
(b) missing pattern II
Figure 5: Histogram of estimated model parameter $\phi_{1}$ when the data are generated from Model 4 (I(1)). The blue line is the smoothed density estimated from the histogram of the parameter estimated from the ground truth.
In conclusion, under the univariate time series models considered in this subsection, which include linear, nonlinear, and nonstationary time series, TWI consistently fares well in all three comparing metrics. It can capture nonlinear dynamics as illustrated by the low Wasserstein error in the imputed marginal distributions, which provides substantial improvements in downstream ACF and model estimates.
5.2 Multivariate time series
In this subsection, we consider the following two multivariate nonlinear time series models.
Model 6 (Nonlinear VAR; NLVAR).
$x_{t,1}=$ $\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}$
$x_{t,2}=$ $\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
Model 7 (Additive logistic model; AL, Brunsdon and Smith,, 1998).
$\begin{pmatrix}y_{t,1}\\ y_{t,2}\end{pmatrix}=\begin{pmatrix}0.1\\ 0.1\end{pmatrix}+\begin{pmatrix}0.7&-0.5\\ 0&-0.7\end{pmatrix}\begin{pmatrix}y_{t-1,1}\\ y_{t-1,2}\end{pmatrix}+0.2\begin{pmatrix}\epsilon_{t,1}\\ \epsilon_{t,2}\end{pmatrix},$
@@ 5.2 Multivariate time series
Model 6 is a VAR model with a nonlinear term. In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function. Model 7 generates the so-called compositional time series, which refers to a multivariate time series $\{\mathbf{x}_{t}\}$, where $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$, $d>1$, such that for each $t$, $\sum_{j=1}^{d}x_{t,j}=1$, and $x_{t,j}\geq 0$ for all $j$. Such data are encountered in a wide array of fields such as the environmental sciences, economics, geology and political science. Here, we adopt the additive logistic (AL) model of Brunsdon and Smith, (1998) in Model 7. The AL model postulates that after some nonlinear transformation, the data follow a standard VARMA model. In fact, one can view compositional data as a time series on some low-dimensional manifold. Alternative models for composition time series include the recent works of Zheng and Chen, (2017), Harvey et al., (2024), and Zhu and Müller, (2024).
Focus on Model 6 (NLVAR) first. The Wasserstein distances between the 3-dimensional marginal distributions of the imputed and the original series, defined in (18), are again reported Table 1. TWI methods clearly produce the lowest Wasserstein distances. As depicted in Figure 6, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships. Next, we assess the effectiveness of the imputed series in estimating the model parameters, assuming the nonlinearity is known. That is, we estimate $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$ by plugging in the imputed series. The root mean square estimation errors are documented in Table 3. Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$. It attains more than 50% reduction in RMSE from initialization under both missing patterns I and II, and can achieve 50% lower RMSE than the scalar filter, the best-performing benchmark.
(a) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,1}$
(b) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,1}$
(c) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,2}$
(d) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,2}$
Figure 6: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 (NLVAR) under missing pattern II.
Next, we turn to Model 7 (AL). Since the components must sum to unity, they exhibit negative correlation when one component approaches the endpoints 0 and 1. However, such correlation tends to weaken when the series are near the midpoint. Table 1 records the Wasserstein distance between the 3-dimensional marginal distribution of the imputed and original series. With the smallest Wasserstein loss, TWI methods again demonstrate their ability to adapt to nonlinear dependence structures.
@@
It is worth noting that while the linear interpolation results in a compositional series, other methods, such as the Kalman smoothing, do not generally maintain this compositional property. By restricting the admissible set, TWI ensures that the imputed series remain compositional. As shown in Figure 7, linear interpolation falsely promotes negative correlation between $\hat{w}_{t,3}$ and $\hat{w}_{t-1,2}$ and Kalman smoothing often deviates from the plausible range of a compositional series. TWI amends these peculiarities, leading to imputations that align more closely with the underlying data.
@@
Figure 7: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data (plotted in blue) and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series (plotted in red) for $i,j=1,2,3$, $i\neq j$, when the data are generated from Model 7 (AL) under missing pattern I.
6 Groundwater data application
In this section, we apply the proposed temporal Wasserstein imputation to a groundwater dataset from Taiwan. Groundwater serves as a critical water source for Taiwan, and its management is of considerable importance to the policymakers as well as the industrial sector, particularly due to the island’s steep topography and varying rainfall distribution (Hsu et al.,, 2020). However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells. To address this, imputation of missing values is essential to facilitate reliable downstream analyses.
6.1 Preprocessing and exploratory analysis
Our goal is to study the seasonal groundwater variations, for which we first transform the data into monthly records. The data contain hourly observations from 352 monitoring sites between October, 1992 and August, 2020. For each site, we define monthly groundwater level by the average of the observations in a given month. If more than one-third of the observations for a month are missing, we classify that month as missing. For our analysis, we focus on sites with a less-than-median amount of missing data, reducing the sample to 176 sites. Figure 8 shows the monthly groundwater levels for these sites, after subtracting the mean and normalized by the standard deviation. We make two observations. First, many series exhibit strong seasonality, which is due to the rainfall pattern in Taiwan. There is more rain in the summer due to monsoon and typhoons, while winters are generally drier. Second, there are many outlying values and anomalies in the observed series. For instance, there are abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean. For the purpose of our analysis, we treat such extreme values as missing data.
During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites. Figure 9 shows the fluctuations in the percentage of sites with missing data over time. Although missing data are mostly prevalent in the early years with more than 50% of the sites generating missing data, there are some time during which the percentage surges, such as between 2003 and 2005. In addition, the percentage of missing data is gradually increasing in more recent years. Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
Figure 9: Monthly percentage of sites recording a missing value.
6.2 Results
Similar to Section 5, we employ linear interpolation, spline interpolation, Kalman smoothing, and the scalar filter as benchmark imputation methods. For TWI and $k$-TWI, linear interpolation and Kalman smoothing are again used as initializations.
We consider two downstream tasks using the imputed series. First, we perform dimension reduction via the time series factor model of Lam et al., (2011). The factor model postulates that the observations $\mathbf{x}_{t}\in\mathbb{R}^{176}$ are driven by a small number of latent factor processes. Specifically,
$\mathbf{x}_{t}=\mathbf{A}\mathbf{f}_{t}+\bm{\epsilon}_{t},$
where $\mathbf{f}_{t}\in\mathbb{R}^{r}$, $r\ll 176$, is the factor process and $\{\bm{\epsilon}_{t}\}$ is white noise. Factor modeling is helpful for understanding the high-dimensional time series and its complex temporal patterns. However, estimating the factor model requires eigenanalysis of the autocovariance matrices $\mathbf{\Sigma}(h)=\mathbb{E}(\mathbf{x}_{t}\mathbf{x}_{t-h})$, which becomes infeasible when the series are severely corrupted by missing data.
Figure 10 shows the leading two factors estimated using series imputed by the benchmark methods. The linear and spline interpolation methods, both favoring trend preservation, result in the first estimated factor displaying strong trends, while the second factor reflects more the seasonal patterns. Notably, this is reversed when factors are estimated based on series imputed by Kalman smoothing and scalar filter. In these cases, the leading factor process exhibits strong seasonality, and the second factor displays trend patterns with a slight U-shape, which is quite different from those estimated with linear and spline methods.
Figure 10: Factor processes estimated using series imputed by the benchmark methods. The dashed red line marks January 2002.
Figure 11 displays the leading two factors estimated using series imputed by TWI and $k$-TWI. Note that while linear interpolation has led to a strong trend factor, as shown in Figure 10, the leading estimated factor of TWI initialized by linear interpolations exhibits significant seasonal pattern, which aligns more closely with those estimated by Kalman smoothing and scalar filter. Moreover, the second estimated factor also shows trends resembling those of the Kalman smoothing and scalar filter, but with a more apparent U-shape. To compare TWI with Kalman smoothing, we focus on the data before January 2002 (marked by the red dashed lines in Figure 10 and 11). Recall that missing data are prevalent during this period (Figure 9). Factors estimated by Kalman smoothing are affected by the abundance of missing data and tend to show smaller variations in the early periods, which may be the result of small variances in the optimal interpolator. In contrast, factors estimated using data imputed by TWI (initialized by Kalman smoothing) show consistent variational patterns across time and appear more trustworthy in understanding the dynamics. Overall, TWI took a more holistic view of the series which accounts for the prominent seasonality and trend in the data.
Figure 11: Factor processes estimated using series imputed by TWI. The dashed red line marks January 2002.
@@ 6.2 Results
Our second downstream task concerns autoregreesive modeling. Specifically, we fit an AR(6) model using data imputed by the methods considered. The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including missing values. Then, we assess the impact of imputation quality by the out-of-sample prediction performance in predicting the remaining data. Note that the missing data in the test data are removed for fair comparison. Figure 12 presents, in violin plots, the distribution of the root mean squared prediction errors (RMSE) for each method across the 176 series under study. The TWI-type methods effectively alleviate the large RMSEs caused by the benchmark imputation methods and produce lower RMSEs in general.
@@
Figure 12: The root mean squared prediction errors (RMSEs) distribution for each method across the 176 series, ranked by median from left to right.
In conclusion, the temporal Wasserstein imputation preserves important temporal features such as the seasonal patterns and enables complex downstream statistical analyses such as factor modeling. Our analysis show that the results obtained from TWI appear reliable. Thus, we expect the series imputed by TWI are also suitable for further analyses, especially when nonlinear models are employed.
7 Conclusion
In this work, we proposed an imputation method, the temporal Wasserstein imputation, for imputing missing entries in time series data. The proposed method complements the existing approaches by effectively capturing nonlinear dynamics and reducing the distributional bias typically observed for the conventional methods in handling missing time series data. Hence, the imputation of the proposed method is more reliable in downstream analysis. Our simulation studies and the analysis of groundwater data support the practical utility of the proposed method.
Appendix A Proof of Proposition 6
Lemma 8.
Suppose $\{\tilde{x}_{t}\}$ satisfies (A1) and (A2). Then
where $\tilde{\mathbf{x}}=(\tilde{x}_{0},\ldots,\tilde{x}_{n-1})$.
Proof.
$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}\Rightarrow\mu_{p},$ (20)
and
$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|^{k}.$ (21)
By Theorem 36.4 of Billingsley, (1995), the processes $\{\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\}_{t}$ and $\{\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\}_{t}$ are stationary and ergodic for any $\mathbf{h}\in\mathbb{R}^{p}$, where $\mathrm{i}=\sqrt{-1}$. Thus (21) follows immediately from the Ergodic theorem. To show (20), it suffices to show, with probability one,
By the Ergodic theorem, we have
$\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q}),\quad\mbox{for all }\mathbf{q}\in\mathbb{Q}^{p},$ (22)
with probability one. Let
noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2). By the Ergodic theorem and (22), $\mathbb{P}(\mathcal{E})=1$. Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary. Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
$|\varphi(\mathbf{h})-\varphi(\mathbf{q})|<\epsilon,$ (23)
$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (24)
for all $\mathbf{h}$, $\mathbf{q}$ such that $\|\mathbf{h}-\mathbf{q}\|<\delta$. Combining (23)–(24), we have
on $\omega\in\mathcal{E}$. Note, in particular, that $\epsilon$ and $\delta$ do not depend on $\mathbf{h}$. Letting $\epsilon\rightarrow 0$ and $\delta\rightarrow 0$ proves (20) holds on $\mathcal{E}$, and hence with probability one. ∎
Proof of Proposition 6.
Now the desired result follows from Lemma 8. ∎
Appendix B Scalar filter algorithm
    $\hat{w}_{s}=\hat{w}_{s}^{(0)}-\hat{\beta}^{(s)},\quad s\in\mathcal{M}.$

### v2

@@ 5.2 Multivariate time series
Figure 3: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 under missing pattern II.
6 Groundwater data application
In this section, we apply the proposed TWI to a groundwater dataset from Taiwan, which was collected and analyzed by Hsu et al., (2020). Given the island’s steep topography and varying rainfall distribution, groundwater management is of considerable importance for Taiwan, particularly as it is a critical water source. However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells.
We aim to study the seasonal groundwater variations, for which we first transform the hourly data from 352 sites between October 1992 and August 2020 into monthly records. For each site, monthly groundwater level is calculated as the average of the observations in a given month, with months classified as missing if more than one-third of the observations are missing. Our analysis focuses on sites with a less-than-median amount of missing data, reducing the sample to 176 sites. Figure 4 shows the monthly groundwater levels for these sites, after subtracting the mean and normalizing by the standard deviation. We make two observations. First, many series exhibit strong seasonality due to the rainfall pattern in Taiwan. Monsoon and typhoons increase rainfall in summer, while winters are drier. Second, there are many outlying values and anomalies in the observed series, including abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean. We also treat such extreme values as missing data.
During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites. Figure 5 shows the fluctuations in the percentage of sites with missing data over time. Although missing data are mostly prevalent in the early years, there are some time during which the percentage surges, such as between 2003 and 2005. In addition, the percentage is gradually increasing in more recent years. Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
We consider two downstream tasks using the imputed series. First, following Hsu et al., (2020), we decompose the de-trended time series using the independent component analysis (ICA), for which we apply the fastICA algorithm of Hyvärinen and Oja, (2000). Because direct application of ICA becomes infeasible when the series are severely corrupted by missing data, the analysis of Hsu et al., (2020) is confined to a relatively short time span. The second downstream task considers autoregressive modeling, which is presented in Section S3.3 in the supplementary material. In implementation, we chose $n_{1}=160$ and $p=6$ for TWI, and for $k$-TWI we use $n_{1}\in\{160,225,110\}$.
Figure 5: Monthly percentage of sites recording a missing value.
Figure 6 shows the leading two independent components (ICs) estimated using the imputed series. ICs estimated from series imputed by Kalman smoothing have peaks around August-September (red shaded area) and May-June (blue shaded area), which was also observed in Hsu et al., (2020). These peaks are attributed to the dual rain seasons in Taiwan, which involve typhoons in August-September and the Meiyu period from mid-May to mid-June. ICs estimated from TWI-imputed series with Kalman initialization largely preserve these seasonal peaks, but with improved performance before 2002. Recall that missing data are prevalent during this period (Figure 5). Especially for the first IC, those computed from TWI-imputed series have clear peaks around August-September with a more consistent amplitude. On the other hand, the ICs computed from the linearly interpolated series could not effectively isolate and identify the Meiyu season, with peaks occurring slightly earlier than expected. Nonetheless, ICs estimated from TWI-imputed series with linear interpolation substantially alleviates this shortcoming and produced more interpretable results. ICs computed from spline-interpolated series also identifies the two rain seasons, albeit they show more rugged trajectories instead of clean periodic patterns. For the scalar filter, the resulting ICs similarly appear noisy, and the Meiyu season is not well-isolated, showing that the imputation induces distortions in the dynamic pattern.
Figure 6: Independent components estimated using the imputed series. The dashed red lines in each plot mark January 2002. The red and blue shaded areas correspond to August-September (typhoon season) and May-June (Meiyu season), respectively.
Overall, TWI took a more holistic view which accounts for the underlying dynamics, such as complex seasonality, and produced interpretable results. In Section S3.3 of the supplementary material, we conduct further experiments to show that the marginal distributions of TWI-imputed series, in terms of AR modeling, are also more consistent across time compared to other methods. These results show TWI performs reliably in practice and is a useful method in handling missing data in time series.
SUPPLEMENTARY MATERIAL
Supplementary material:
The file contains the proofs of the propositions presented in Section 4 and some technical results. It also documents the scalar filter algorithm of Peña and Tsay, (2021) and complementary numerical results, including the simulation results for autocovariance estimation and for data generated from a compositional time series model. Finally, it includes the analysis of the groundwater data via autoregressive modeling.
Supplementary material to “Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series”
This supplementary material contains three sections. Section S1 collects the proofs of the propositions presented in main text as well as some technical results. Section S2 documents the scalar filter algorithm of Peña and Tsay, (2021) for completeness. Section S3 contains complementary numerical results, including the simulation results for autocovariance estimation and imputing data generated from a compositional time series model. It also includes the analysis of autoregressive modeling using the imputed groundwater data.
S1 Technical proofs
Proof of Proposition 3.
Moreover,
$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}^{2}}=$ $\mathrm{diag}\left(\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):i=p-1,\ldots,n_{1}\right),$ (S1.1)
$\frac{\partial^{2}F_{0}}{\partial\mathbf{v}^{2}}=$ $\mathrm{diag}\left(\sum_{i=p-1}^{n_{1}}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):j=n_{1}+1,\ldots,n-1\right),$ (S1.2)
$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}\partial\mathbf{v}}=$ $-\left(\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j})\right)_{p-1\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}.$ (S1.3)
Since $\ell$ is convex, ′′$\ell^{\prime\prime}\geq 0$. From (S1.1)–(S1.3), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries. It follows from the Gershgorin circle theorem that $\mathbf{H}$ is positive semi-definite. Thus $F_{0}$ is convex in $\mathbf{w}$. ∎
Before proving Proposition 4, we first state our assumptions and establish Lemma S6.
The underlying time series $\{\tilde{x}_{t}\}$ is stationary and ergodic.
$\mathbb{E}|\tilde{x}_{t}|^{k}<\infty$ for some $k\geq 1$.
These assumptions are satisfied by many linear and nonlinear time series models.
Lemma S6.
Suppose $\{\tilde{x}_{t}\}$ satisfies (A1) and (A2). Then
where $\tilde{\mathbf{x}}=(\tilde{x}_{0},\ldots,\tilde{x}_{n-1})$.
Proof.
$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}\Rightarrow\mu_{p},$ (S1.4)
and
$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|^{k}.$ (S1.5)
By Theorem 36.4 of Billingsley, (1995), the processes $\{\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\}_{t}$ and $\{\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\}_{t}$ are stationary and ergodic for any $\mathbf{h}\in\mathbb{R}^{p}$, where $\mathrm{i}=\sqrt{-1}$. Thus (S1.5) follows immediately from the Ergodic theorem. To show (S1.4), it suffices to show, with probability one,
By the Ergodic theorem, we have
$\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q}),\quad\mbox{for all }\mathbf{q}\in\mathbb{Q}^{p},$ (S1.6)
with probability one. Let
noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2). By the Ergodic theorem and (S1.6), $\mathbb{P}(\mathcal{E})=1$. Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary. Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
$|\varphi(\mathbf{h})-\varphi(\mathbf{q})|<\epsilon,$ (S1.7)
$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (S1.8)
for all $\mathbf{h}$, $\mathbf{q}$ such that $\|\mathbf{h}-\mathbf{q}\|<\delta$. Combining (S1.7)–(S1.8), we have
on $\omega\in\mathcal{E}$. Note, in particular, that $\epsilon$ and $\delta$ do not depend on $\mathbf{h}$. Letting $\epsilon\rightarrow 0$ and $\delta\rightarrow 0$ proves (S1.4) holds on $\mathcal{E}$, and hence with probability one. ∎
Proof of Proposition 4.
Now the desired result follows from Lemma S6. ∎
S2 Scalar filter algorithm
    $\hat{w}_{s}=\hat{w}_{s}^{(0)}-\hat{\beta}^{(s)},\quad s\in\mathcal{M}.$
S3 Complementary numerical results
S3.1 Autocovariance estimation
In this subsection, we compare the ACFs estimated from the imputed series. Define $\gamma(\mathbf{z},h)=(n-h)^{-1}\sum_{t=h}^{n-1}(z_{t}-\bar{z})(z_{t-h}-\bar{z})$, where $\mathbf{z}=(z_{0},z_{1},\ldots,z_{n-1})$ and $\bar{z}=n^{-1}\sum_{t=0}^{n-1}z_{t}$. Then the RMSE is defined as
$\mathrm{RMSE}_{\gamma}(h)=\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}(\gamma(\hat{\mathbf{w}}^{(i)},h)-\gamma_{\mathbf{x}}(h))^{2}},$ (S3.9)
where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(x_{0}^{(i)},\ldots,x_{n-1}^{(i)})$.
@@ 6 Groundwater data application
Table S1 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$. For Model 1 (AR), most methods perform reasonably well. Linear interpolation produces accurate ACF estimates since the positive AR coefficient intrinsically favors linearity, while the scalar filter exhibits largest estimation errors. For Model 2 (ARMA), TWI improves from linear and Kalman initializations and yields more accurate ACF estimates: TWI reduces the biases in Kalman smoothing when $h=0$ and in linear interpolation when $h=1$, and $h=2$. For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series exhibit the lowest RMSEs, with substantial improvements over the initializations. Notably, the scalar filter also yields accurate ACFs for $h=1,2$. Figure S1 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
@@
Figure S1: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR). The blue line is the smoothed density estimated from the histogram of the ACFs estimated from the ground truth.
Table S1: $\mathrm{RMSE}_{\gamma}(h)$, defined in (S3.9), for $h=0,1,2$, with $h$ denoting lag. The results are averaged over 1000 simulations, and the smallest values in each row are marked in boldface. See Table 1 in the main text for notations.
@@
S3.2 Compositional time series
In this subsection, we discuss the simulation results associated with the following model.
DGP 7 (Additive logistic model; AL, Brunsdon and Smith,, 1998).
$y_{t,1}=$ $0.1+0.7y_{t-1,1}-0.5y_{t-1,2}+0.2\epsilon_{t,1}$
$y_{t,2}=$ $0.1-0.7y_{t-1,2}+0.2\epsilon_{t,2}$
@@
DGP 7 generates the so-called compositional time series, which refers to a multivariate time series $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$, $d>1$, such that for each $t$, $\sum_{j=1}^{d}x_{t,j}=1$, and $x_{t,j}\geq 0$ for all $j$. Such data are encountered in a wide array of fields such as the environmental sciences, economics, geology and political science. Here, we adopt the additive logistic (AL) model of Brunsdon and Smith, (1998) in DGP 7. The AL model postulates that after some nonlinear transformation, the data follow a standard VARMA model. Alternative models for composition time series include the recent works of Zheng and Chen, (2017), Harvey et al., (2024), and Zhu and Müller, (2024).
@@
Table 1 in the main paper records the Wasserstein distance between the 3-dimensional marginal distributions of the imputed and original series. With the smallest Wasserstein loss, TWI methods again demonstrate their ability to adapt to nonlinear dependence structures. It is worth noting that while the linear interpolation results in a compositional series, other methods, such as the Kalman smoothing, do not generally maintain this compositional property. By restricting the admissible set, TWI ensures that the imputed series remain compositional. As shown in Figure S2, linear interpolation falsely promotes negative correlation between $\hat{w}_{t,3}$ and $\hat{w}_{t-1,2}$ and Kalman smoothing often deviates from the plausible range of a compositional series. TWI amends these peculiarities, leading to imputations that align more closely with the underlying data.
@@
Figure S2: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data (plotted in blue) and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series (plotted in red) for $i,j=1,2,3$, $i\neq j$, when the data are generated from Model 7 (AL) under missing pattern I.
S3.3 Autoregressive modeling with groundwater data
@@
As our second downstream task for the groundwater data analysis, we fit an AR(6) model using data imputed by the methods considered. The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including imputed values. Then, we assess the impact of imputation quality by the out-of-sample prediction performance in predicting the remaining data. The missing data in the test data are removed for fair comparison. Figure S3 presents, in violin plots, the distribution of the logarithm of the root mean squared prediction errors ($\log$ RMSE) for each method across the 176 series under study. The TWI-type methods effectively alleviate some large RMSEs caused by the benchmark imputation methods and generally produce lower RMSEs. This shows the series imputed by TWI have marginal distributions that are stable (across time), meaning the method does not artificially induce structural shifts. This further corroborates the practical utility of TWI.
@@
Figure S3: The distributions of the logarithm of the root mean squared prediction errors (RMSEs) for each method across the 176 series, ranked by median from left to right.